OpenStudy (anonymous):

If y is 6 when x is 3 and y varies directly with x, find y when x is 8. 4 11 16 i dont understand this its worded wered

OpenStudy (anonymous):

Well in what is given to you, if y varies directly with x than your equation is: y/x=k. When you plug your given numbers in you will get:6/3=k and when solving for k you recieve k=2. Now in the wording of this (it looks like an algebra 1 student created this sheet) when it says "find y when x is 8" i would guess that means that the y variables are the three following numbers you listed (4,11,16) and the x is of course 8. Just assume they're all direct variation and use the formula y/x=k to solve for them.

OpenStudy (anonymous):

so 16

OpenStudy (anonymous):

you're going to have to be more clear i have no idea how or what you got 'so 16' from

OpenStudy (anonymous):

the awnser

OpenStudy (anonymous):

I think you need a tudor.

OpenStudy (anonymous):

i think you need to shut the fu** up

OpenStudy (anonymous):

Story of my life